# STUPP-24-271

### v1

@@ 3 The Jarlskog invariant and its numerical results
In this section, we investigate the behavior and numerical results of CP violation after imposing the three experimental constraints. The Jarlskog invariant $J$ [10] of the unitary matrix $U_{\rm MNS}$ (18) is
@@
To the extent that $s_{e},s_{13}\lesssim 0.1$ holds, higher orders of $s_{e}$ and $s_{13}$ can be negligible, and the dominant terms are those proportional to $\sin\alpha$ and $\sin\beta$. Note that the relative sign between terms depends only on $s_{e}c_{e}$.
@@
The term with $\sin(\alpha-\beta)$ is a sum of the term with $\sin\alpha$ and a contribution from the expansion of $s_{13}$, while the term with $\cos\alpha$ is from the expansion of $s_{12}$. In the MNS matrix, the second term is dominant over the third term because $S_{23}/S_{13}\simeq 5$ and $(1-2S_{12}^{2})C_{23}/C_{12}S_{12}\simeq 1/2$ hold. The same is true for the CKM matrix from $S_{23}/S_{13}\simeq 10$.
@@ 3.1 Numerical results
Plots of $\sin\delta$ as a function of $\alpha,\beta$ and $s_{e}$ are shown in Fig. 2. The range of $s_{e}$ is set within the critical value $|s_{e}|\leqq 0.2$. The green-shaded region corresponds to $\cos\delta<0$. From Eq. (38), if $s_{e}$ is sufficiently small, $\delta$ is approximately equal to $\beta$. On the other hand, when $s_{e}$ approaches the critical value, it shows singular behavour on the line $\alpha-\beta=\pm\pi/2,\pm 3\pi/2$. For $s_{e}$ above the critical value, the divided two regions correspond to the large and the small solution, respectively.
@@
In particular, for $\beta=0$ or $\pi$, a kind of generalized CP symmetry (GCP) [11, 12, 13, 14, 15] can be defined. In this case, there is an upper bound on $\delta$ from Eq. (38).
$|\sin\delta^{\rm max}|\simeq\left|{s_{e}S_{23}\over S_{13}}\sin\alpha\right|\,.$ (39)
However, this expression is not accurate for $s_{e}\simeq s_{13}$, and a large CPV such as $\delta\simeq 1$ is possible even for $\beta=0$ or $\pi$ and $s_{e}\sim s_{e}^{c}$. On the other hand, the chiral symmetry does not hold for $s_{e}\simeq 0.2$, and the parameter region where $s_{e}$ is close to the critical value is out of scope in this letter.
The long-baseline experiment using HK and J-PARC [3] has a discovery capability of around $\delta\gtrsim 30^{\circ}$ after 10 years of operation. This allows for physical interpretations when HK observes or does not observe the Dirac phase $\delta$, respectively.
When HK observes the Dirac phase (of about $O(1)$)
Large CP violation, as suggested by the T2K experiment, cannot be achieved by the relative phase $\alpha$ alone. Therefore, if the Yukawa matrix of charged leptons $Y_{e}$ is sufficiently hierarchical, large CP violation corresponds to the intrinsic CP violation $\delta_{\nu}$ in the diagonalization $O_{\nu}$ of the neutrino matrix. In particular, if the phase is close to $\pi/2$, $\mu-\tau$ reflection symmetry [16, 17, 18] (and as a review [19]) is more favored.
When HK does not observe the Dirac phase (of about 30 or more)
@@
In such a situation, a kind of CP symmetry such as $\beta=0$ or $\pi$ can be favored. However, even for large $s_{e}\sim 0.1$, the relative phase $\sin\alpha$ is hardly restricted (and $\alpha\sim\pm\pi/2$ is allowed). The smaller the observed CPV, the more important the relative phase can be for the source of CPV. Such restrictions on the CP phase can provide important constraints on the leptogenesis scenario [20].
@@ 3.2 CKM matrix
A similar analysis is performed for the CKM matrix. As input parameters, we use the latest data of UTfit [21].
$\sin\theta_{12}$ $=0.22519\pm 0.00083\,,~{}~{}~{}\sin\theta_{23}=0.04200\pm 0.00047\,,$ (41)
$\sin\theta_{13}$ $=0.003714\pm 0.000092\,,~{}~{}~{}\delta_{\rm KM}=1.137\pm 0.022\,.$ (42)
The parameter $s_{e}$ in the MNS matrix is redefined as $s_{u}$ in this case, and a similar analysis for the CKM matrix is shown in Fig. 3. The range of $s_{u}$ is set to be within the critical value ($|s_{u}|\leqq 0.09$) and the region within the error range of $\delta_{\rm KM}$ was filled in. The green-shaded regions are excluded because $\cos\delta_{\rm KM}<0$. In the three continuous parameters, one two-dimensional surface is chosen that reproduces the observed CP phase.
@@
Since the phase behaves as $\delta_{\rm KM}\simeq\beta$ for sufficiently small $s_{u}$, we already see that the observed CPV in the CKM matrix corresponds to the intrinsic CP phase associated with the diagonalization of down quarks. On the other hand, near the critical value of $s_{u}\simeq 0.09$, a parameter set $\beta=0$ or $\pi$ and $\alpha\simeq\pm\pi/2$ is possible. In the limit of $s_{13}\to 0$, Eq. (18) is reduced to the Fritzsch–Xing parameterization and it has been known that the phase is almost maximal $\alpha=\pi/2$ [8]. This situation is realized by a kind of generalized CP symmetries [22, 23, 24, 25]. Such restrictions can be useful in the construction of grand unified theories.
@@ 4 Summary
In this letter, we perform an almost general analysis of flavor mixing matrices to investigate the discriminative power of CP phases by the next-generation neutrino oscillation experiments. For hierarchical Yukawa matrice $Y_{f}$ of the Standard Model fermions $f$, the unitary matrix $U_{f}$ of diagonalization for left-handed fields is expected to have small mixings. As an assumption, we neglect the 1-3 mixing of more hierarchical fermions (i.e. up-type quarks and charged leptons). Thus there are two sources of CP violation in the flavor mixing matrices, the intrinsic CP phase $\delta_{d,\nu}$ in diagonalization of less hierarchical fermions $d,\nu$ and relative phases between two unitary matrices.
By eliminating unphysical phases and imposing constraints of the three measured mixing angles, the flavor mixing matrices are analytically displayed by two phases and the 1-2 mixing $s_{u,e}$ of more hierarchical fermions. For sufficiently small 1-2 mixing $s_{e}$ of charged leptons, the Dirac phase $\delta$ is almost identical to the intrinsic phase of neutrinos $\delta_{\nu}$ and contributions of relative phases are not dominant. Therefore, future detection of the Dirac phase by the Hyper-Kamiokande experiment indicates the observation of $\delta_{\nu}$. On the other hand, if such a CP violation is not observed, an upper limit is placed on a combination of $\delta_{\nu}$ and relative phases.

### v2

@@ 3 The Jarlskog invariant and its numerical results
In this section, we investigate numerical behaviors of the CP violation after imposing the three experimental constraints. The Jarlskog invariant $J$ [20] of the unitary matrix $U_{\rm MNS}$ (19) is
@@
To the extent that $s_{e},s_{13}\lesssim 0.1$ holds, higher orders of $s_{e}$ and $s_{13}$ are negligible, and the dominant terms are those proportional to $\sin\alpha$ and $\sin\beta$. Note that the relative sign between terms depends only on $c_{e}s_{e}$.
@@
The term with $\sin(\alpha-\beta)$ is a sum of the term with $\sin\alpha$ and a contribution from the expansion of $s_{13}$, while the term with $\cos\alpha$ is from the expansion of $s_{12}$. For the MNS matrix, the second term is dominant over the third term because $S_{23}/S_{13}\simeq 5$ and $(1-2S_{12}^{2})C_{23}/C_{12}S_{12}\simeq 1/2$ hold. The same is true for the CKM matrix from $S_{23}/S_{13}\simeq 10$.
@@ 3.1 Numerical results
Plots of $\sin\delta$ as a function of $\alpha,\beta$ and $s_{e}$ are shown in Fig. 2. The range of $s_{e}$ is set within the critical value $|s_{e}|\leqq 0.2$. The green-shaded region corresponds to $\cos\delta<0$. From Eq. (39), if $s_{e}$ is sufficiently small, $\delta$ is approximately equal to $\beta\sim\delta_{\nu}$. On the other hand, when $s_{e}$ approaches the critical value $s_{e}^{c}$, it shows singular behavour on the line $\alpha-\beta=\pm\pi/2,\pm 3\pi/2$. For $s_{e}>s_{e}^{c}$, the divided two regions correspond to the large and small solutions in Fig. 1.
@@
The DUNE and T2HK experiment has a discovery capability of around $\delta\gtrsim 30^{\circ}$ after 10 years of operation. This allows for physical interpretations when the next-generation experiments observe or do not observe the Dirac phase $\delta$, respectively.
When the Dirac phase (of about $O(1)$) is observed
Large CP violation, as suggested by the T2K experiment, cannot be achieved by the relative phase $\alpha$ alone. Therefore, if the Yukawa matrix of charged leptons $Y_{e}$ is sufficiently hierarchical, large CP violation corresponds to the intrinsic CP violation $\delta_{\nu}$ in the diagonalization $U_{\nu}^{0}$ of the neutrinos. In particular, if the phase is close to $\pi/2$, the $\mu-\tau$ reflection symmetry [21, 22, 23] (and as a review [24]) is more favored. However, in regions where $|s_{e}\sin\alpha|$ is large, the contribution of $\sin\beta$ can be subdominant.
When the Dirac phase (of about 30 or more) is does not observed
@@
In this situation, $\beta=0$ or $\pi$ can be favored, that is realized by some generalized CP symmetries [25, 26, 27, 28]. The smaller the observed Dirac phase, the more important the relative phase can be for the source of CPV. Besides, even for large $s_{e}\sim 0.1$, the relative phase $\sin\alpha$ is hardly restricted (and $\alpha\sim\pm\pi/2$ is allowed).
The reason for this physical interpretation is as follows: in Eq. (14), the mixing $s_{e}$ and $s_{\tau}$ become zero when the chiral symmetry of the first and second generation becomes exact. In this limit, the phases $\rho$ and $\sigma$ can be removed by further phase redefinition. Therefore, effects of the original phases $\rho,\sigma$ are suppressed by $s_{e}$ and $s_{\tau}$, and $\delta_{\nu}$ becomes the leading CPV source for small $s_{e}$. Such restrictions on CP phases can provide important constraints on the leptogenesis scenario [29].
@@ 3.2 CKM matrix
A similar analysis is performed for the CKM matrix. The 1-3 element of the CKM matrix satisfies
$|V_{ub}|\sim m_{d1}/m_{d3}\gg m_{u1}/m_{u3}\,,$ (41)
and the approximation by chiral symmetry seems to be reasonable. As input parameters, we use the latest data of UTfit [30].
$\sin\theta_{12}^{\rm CKM}$ $=0.22519\pm 0.00083\,,~{}~{}~{}\sin\theta_{23}^{\rm CKM}=0.04200\pm 0.00047\,,$ (42)
$\sin\theta_{13}^{\rm CKM}$ $=0.003714\pm 0.000092\,,~{}~{}~{}\delta_{\rm KM}=1.137\pm 0.022\,.$ (43)
A similar plot for the CKM matrix is shown in Fig. 3. The parameter $s_{e}$ in the MNS matrix is redefined as $s_{u}$ and its range is set to be within the critical value ($|s_{u}|\leqq 0.09$). The green-shaded regions are excluded because $\cos\delta_{\rm KM}<0$. The region within the error range of $\delta_{\rm KM}$ is filled in. In the three continuous parameters, one two-dimensional surface is chosen that reproduces the observed CP phase.
@@
Since the phase behaves as $\delta_{\rm KM}\simeq\beta$ for sufficiently small $s_{u}$, we see that the observed CPV in the CKM matrix corresponds to the intrinsic CP phase $\delta_{d}$ associated with the diagonalization of down quarks. This situation corresponds to the large CPV in the MNS matrix, and $|s_{u}\sin\alpha|$ must be large to make $\sin\beta$ small. A characteristic structure of CP phases such as the $\mu-\tau$ reflection symmetry can be required. On the other hand, near the critical value, a parameter set $s_{u}=0.09,\alpha=-\pi/2$ or $s_{u}=-0.09,\alpha=\pi/2$ is possible. In the limit of $s_{13}\to 0$, Eq. (19) is reduced to the Fritzsch–Xing parameterization [18] and it has been known that the phase is almost maximal $|\alpha|=\pi/2$. Such restrictions can be useful in the construction of grand unified theories.
@@ 4 Summary
In this letter, we perform an almost general analysis of flavor-mixing matrices $V_{\rm CKM}$ and $U_{\rm MNS}$ to investigate the discriminative power of CP phases by next-generation neutrino oscillation experiments. For hierarchical Yukawa matrice $Y_{f}$ of the Standard Model fermions $f$, the unitary matrix $U_{f}$ of diagonalization for left-handed fields is expected to have small mixings. As an approximation, we neglect the 1-3 mixing of more hierarchical fermions (i.e. up-type quarks and charged leptons). Thus there are two sources of CP violation in $V_{\rm CKM}$ and $U_{\rm MNS}$, the intrinsic CP phase $\delta_{d,\nu}$ in diagonalization of less hierarchical fermions $d,\nu$ and relative phases between two unitary matrices.
By eliminating unphysical phases and imposing constraints of the three measured mixing angles, the flavor-mixing matrices are analytically displayed by two phases and the 1-2 mixing $s_{u,e}$ of more hierarchical fermions. For sufficiently small 1-2 mixing $s_{e}$ of charged leptons, the Dirac phase $\delta$ is mostly identical to the intrinsic phase of neutrinos $\delta_{\nu}$ and contributions of relative phases are not dominant. Therefore, future detection of the Dirac phase indicates the observation of $\delta_{\nu}$. On the other hand, if such a CP violation is not observed, an upper limit is placed on a combination of $\delta_{\nu}$ and relative phases. This interpretation holds because the relative phases can be eliminated by further phase redefinition in the limit of zero mixing of the charged leptons. Thus, contributions of relative phases are proportional to mixing angles of the charged leptons, which are suppressed by the chiral symmetries.